# Theory of intervalley-coherent AFM order and topological superconductivity in tWSe2

### v1

@@ Abstract
The recent observation of superconductivity in the vicinity of insulating or Fermi surface reconstructed metallic states has established twisted bilayers of WSe2 as an exciting platform to study the interplay of strong electron-electron interactions, broken symmetries and topology. In this work, we study the emergence of electronic ordering in twisted WSe2 driven by gate-screened Coulomb interactions. Our first-principles treatment begins by constructing moiré Wannier orbitals that faithfully capture the bandstructure and topology of the system and project the gate-screened Coulomb interaction onto them. Using unbiased functional renormalization group calculations, we find an interplay between intervalley-coherent antiferromagnetic order and chiral, mixed-parity $d/p$-wave superconductivity for carrier concentrations near the displacement field-tunable van-Hove singularity. Our microscopic approach establishes incommensurate intervalley-coherent antiferromagnetic spin fluctuations as the dominant electronic mechanism driving the formation of superconductivity in $\theta=5.08^{\circ}$ twisted WSe2 and demonstrates that nesting properties of the Fermi surface sheets near the higher-order van-Hove point cause an asymmetric density dependence of the spin ordering as the density is varied across the van-Hove line, in good agreement with experimental observations. We show how the region of superconducting and magnetic order evolves within the two-dimensional phase space of displacement field and electronic density as twist angle is varied between $4^{\circ}\dots 5^{\circ}$.
@@ I Introduction
Transition metal dichalcogenide (TMD) heterostructures have arisen as fertile and highly versatile platforms for exploring the interplay between topology and strong electronic correlations Kennes et al. (2021), hosting correlated insulators Wang et al. (2020); Xu et al. (2020); Huang et al. (2021), spin-ordered metals Ghiotto et al. (2024); Anderson et al. (2023); Crépel and Fu (2023), rich multi-orbital physics Zhang et al. (2021); Ryee and Wehling (2023), Feschbach resonances Slagle and Fu (2020); Crépel and Fu (2021); Schwartz et al. (2021); Kuhlenkamp et al. (2022); Crépel et al. (2023); Wagner et al. (2023); Lange et al. (2024), quantum critical behavior Ghiotto et al. (2021); Li et al. (2021), excitonic insulators Tran et al. (2019); Ma et al. (2021); Gu et al. (2022); Xiong et al. (2023), topological Kondo physics Dalal and Ruhman (2021); Guerci et al. (2023); Zhao et al. (2023a); Guerci et al. (2024a); Zhao et al. (2023b); Xie et al. (2024a), quantum anomalous and quantum spin-Hall insulators Tao et al. (2024); Kang et al. (2024a), ferromagnetic fractional quantum anomalous phases Zeng et al. (2023); Cai et al. (2023); Xu et al. (2023); Park et al. (2023), or gapped spin liquids Kang et al. (2024b); Crépel and Millis (2024a). Additionally, early experiments on $5.0^{\circ}$ twisted WSe2 hinted at the existence of superconducting phases in moiré TMDs Wang et al. (2020); a finding recently confirmed by sub-Kelvin transport measurements on $3.65^{\circ}$ Xia et al. (2024) and $5.1^{\circ}$ Guo et al. (2024) twisted samples. The discovery of superconductivity in these highly tunable moiré TMDs has sparked many theoretical studies Schrade and Fu (2021); Wietek et al. (2022); Klebl et al. (2023); Kim et al. (2024); Zhu et al. (2024); Christos et al. (2024); Guerci et al. (2024b); Tuo et al. (2024); Chubukov and Varma (2024); Xie et al. (2024b); Qin et al. (2024); Chubukov and Varma (2024) for the new insights it promises to provide on the origin and properties of strongly correlated topological superconductivity, including exciting prospects for trapping non-abelian anyons at clean, gate-defined interfaces Mong et al. (2014); Vaezi (2014); Barkeshli (2016); Crépel and Regnault (2024).
Twisted WSe2 (tWSe2) consists of two WSe2 layers stacked at a relative twist angle $\theta$ which are encapsulated in a structure including top and bottom gates. The twist, in combination with the electrostatic potential from the local arrangement of WSe2 bilayers and the interlayer electronic hybridization, produces a large unit cell with two dimensional “moiré” bands Wu et al. (2019); Devakul et al. (2021). The average of the potentials on the top and bottom gates tunes the electronic density over wide ranges. The difference in voltage between the top and bottom gates introduces an interlayer potential difference, or “displacement field”, which shifts electronic wave functions and charge density between the layers, modifying the electronic band structure. At a fixed twist angle, the system’s behavior can be explored in a phase diagram defined by electronic density and displacement field. As the density varies at a fixed displacement field, the Fermi surface can pass through a van-Hove singularity (VHS), where the density of states diverges. Of particular relevance for this study is that both the density at which a van-Hove singularity occurs and the strength of the resulting divergence in the density of states depend on twist angle and are tunable over wide ranges by varying the displacement field.
In the $5.1^{\circ}$ tWSe2 device Guo et al. (2024), the superconducting state is observed only for a small range of displacement fields near the hole density $n/n_{0}$ at which the Fermi surface passes through the van-Hove singularity, and is adjacent to a metallic region characterized by an increase in longitudinal resistance and a very rapid non-monotonic doping dependence of the Hall resistivity. This unusual metallic region has been interpreted Guo et al. (2024) as arising from the emergence of antiferromagnetic (AFM) spin ordering that partially gaps out the Fermi surface. The densities at which superconductivity and antiferromagnetism occur are not close to one hole per moiré unit cell and the experimentally important displacement fields are significant, on the order of half of the field required to fully layer-polarize the doped holes. In the $3.65^{\circ}$ device Xia et al. (2024), superconductivity is found only for carrier concentrations near $n/n_{0}=1$ per moiré unit cell as well as small displacement fields, with the superconductivity giving way to an insulating phase and then to an apparently conventional metallic phase as the displacement field is increased. In each device, the superconducting and AFM or insulating states are found only in a relatively small region of the experimental phase space defined by carrier concentration and interlayer (displacement field) potential difference.
@@
The association of superconductivity with proximity to the non-integer carrier concentration at which the Fermi surface contains the VHS, and the proximity to an apparent metallic antiferromagnetic phase suggests a weak-to-moderate coupling scenario in the $5.1^{\circ}$ device, where the effects of electron-electron interactions are amplified by the density of states divergence associated with the van-Hove point. The association of superconductivity at particular values of the displacement field may then be related to proximity to a higher order van-Hove singularity (HOVHS) with a stronger density of states divergence appearing as the band theory is tuned Bi and Fu (2021). Recent literature has presented various versions of this scenario. Interactions considered include spin-fluctuations Klebl et al. (2023); Schrade and Fu (2021); Chubukov and Varma (2024), over-screened Coulomb interactions Guerci et al. (2024b) and phenomenological attraction Zhu et al. (2024); Tuo et al. (2024). While these treatments provide important insights, they are based on particular ansätze and come to different conclusions regarding the symmetry of the superconducting state.
@@
Moreover, theory should take into account the non-trivial topological properties Devakul et al. (2021); Crépel and Millis (2024b); Zhang et al. (2024a) of the low-lying hole bands. In addition to potentially endowing the superconducting state with interesting properties Guerci et al. (2024b), they obstruct an effective single-orbital description of the topmost moiré valence band and imply that interactions projected onto the top-most band generically have a long-ranged structure. A minimal faithful low-energy model of tWSe2 comprises three orbitals per valley that are centered on the triangular and honeycomb sites of the moiré superlattice Crépel and Millis (2024b); Qiu et al. (2023). Such a model accurately accounts for the topological entanglement of the topmost valence bands with Bloch states far below the Fermi energy and for the subtle aspects of the fermiology and its dependence on the displacement field. In this multi-orbital Wannier basis, the gate-screened Coulomb interactions relevant for the experimental setup can be cast into fairly local interactions, which ensures the applicability of established many-body techniques beyond the mean-field level Tuo et al. (2024); Guerci et al. (2024b), including the functional renormalization group Salmhofer and Honerkamp (2001); Metzner et al. (2012); Platt et al. (2013); Dupuis et al. (2021); Beyer et al. (2022) (FRG), dynamical mean-field theory Georges et al. (1996), or tensor network based approaches Schollwöck (2011); Orús (2014). A similar situation arises in twisted bilayers of graphene, where a multi-band “heavy fermion” description is argued to best represent the physics Carr et al. (2019a); Song and Bernevig (2022).
@@
In this work, we go beyond initial phenomenological treatments of correlated phases in tWSe2 and provide a full-fledged analysis of correlated states in the weak-to-moderate interaction regime based on physically realistic repulsive gate-screened Coulomb interactions projected onto a multiband model based on exponentially localized Wannier functions that captures the topology and the intricacies of the band structure. Using unbiased functional renormalization group techniques, we map out the interacting phase diagram of tWSe2 in the twist angle range $\theta=4^{\circ}\dots 5^{\circ}$ from first principles. We demonstrate that intervalley-coherent anti-ferromagnetic (IVC-AFM) order and extended regions of chiral $d/p$-wave superconductivity coexist along the gate-tunable van-Hove line. The unbiased characterization of particle-hole and particle-particle instabilities in conjuction with the subtle fermiology near the van-Hove line enabled by our choice of methods permits us to show that the electronic driving mechanism for superconductivity in tWSe2 is primarily facilitated by spin fluctuations around regions of incommensurate IVC-AFM order below the higher order van-Hove singularity, where the absence of perfect nesting strengthens particle-particle fluctuations. Our microscopic approach can explain several features observed in recent experimental measurements, including the unusual density dependence of the particle-hole instabilities with respect to the van-Hove line. We predict a successive shift of superconducting regions to lower values of the displacement field and towards $n/n_{0}\sim 1$ as the twist angle is decreased and briefly discuss the connection of our results to observations on the $\theta=3.65^{\circ}$ sample, where superconductivity is observed for $n/n_{0}\sim 1$ and for small displacement fields.
@@
Figure 1: Fermiology and Wannierization of $\theta=5.08^{\circ}$ tWSe2 for an external displacement field $E_{z}=20\,\mathrm{meV}$. (a) Continuum bandstructure in valley $\nu=+1$ (black) and spectral overlap of its Bloch states with a three-orbital model comprising Wannier orbitals centered at the triangular ($T$, $s$-orbital) and honeycomb ($H_{1}$, $p_{+}$-orbital)/($H_{2}$, $p_{+}$-orbital) sites of the moiré superlattice (tri-colored map). The bandstructure of the effective Wannier model truncated after 4.5 moiré lattice vectors is indicated by the green dashed line. (b) Amplitude $|\psi_{l}\bm{(}r)|^{2}$ of the Wannier functions ($T,H_{1},H_{2}$) on the two layers ($l=0,1$) that capture the spectral weight of the continuum model. (c) Density of states as function of holes per moiré unit cell $n/n_{0}$ and displacement field $E_{z}$. The upper panel shows snapshots of the Fermi surface along the van-Hove line and its orbital polarization using the same colormap as in (a) for valley $\nu=+1$. The Fermi surface of the opposite valley $\nu=-1$ is shown in gray. The green dashed line indicates the mini-Brillouin zone of tWSe2. The blue circle indicates the position of the higher-order VHS, where three van-Hove points merge at $K^{\nu}$ as demonstrated in the insets. The three van-Hove points are moved towards $\Gamma$ ($M$) for increasing (decreasing) value of the external displacement field along the van-Hove line as indicated by the blue triangle (square).
II Model and method
II.1 Multi-orbital Wannier model
@@
The non-interacting electronic structure of hole-doped tWSe2 is accurately described by a continuum model in which the spin-valley locked holes of both monolayers are treated within the effective mass approximation, and are coupled by moiré potentials that slowly vary on the atomic scale Wu et al. (2019). Due to the emergent spin-valley locking in monolayer WSe2, the low-energy states can be assigned a well-defined valley index $\nu=\pm 1$ and states from opposite spin-valley sectors are related by virtue of time-reversal symmetry $\mathcal{T}$. The effects of out-of-plane electric fields are captured by a layer-dependent potential energy ±$\pm E_{z}$. For twist angles $\theta\gtrsim 3^{\circ}$, in-plane lattice relaxation is weak, and the associated moiré fields are characterized by only a few harmonics Carr et al. (2018); Zhang et al. (2024b); Jia et al. (2024); Thompson et al. (2024); Liu et al. (2024). Here, we use the continuum model from Ref. Devakul et al. (2021), whose coefficients have been fitted to large-scale ab-initio calculations, see Supplementary Materials (SM) SM for details.
@@
As most many-body methods require a symmetry-preserving lattice description of the system under scrutiny, we seek to construct a Wannier representation of the topmost valence band states in each valley based on the three-orbital model put forward in Refs. Crépel and Millis (2024b); Qiu et al. (2023). This model is based on generic symmetry arguments and comprises one triangular lattice site $T$ (at Wyckoff position 1a of the triangular superlattice), which has weight on both layers, as well as two honeycomb lattice sites $H_{1,2}$ (at Wyckoff positions 2a,b) with dominant weight on either one of the layers. At a twist angle of $\theta=5.08^{\circ}$, however, the Chern number sequence of the first three valence bands of tWSe2 in valley $\nu=\pm 1$ is $(\nu,\nu,\nu)$ for the parameters of Ref. Devakul et al. (2021), prohibiting an exponential localization of Wannier functions centered at aforementioned Wyckoff positions constructed from the three topmost bands alone. We resolve this issue by employing a single-shot Wannierization scheme Marzari et al. (2012); Carr et al. (2019b); Fischer et al. (2024) in which Bloch states from the entire valence band manifold are considered for the projection and a proper subspace selection is instead achieved by weighting Bloch states according to their energetic hierarchy and their weight around the desired orbital center. Compared to the layer-polarization scheme presented in Ref. Crépel and Millis (2024b), the energetic subspace selection ensures to construct exponentially localized Wannier functions irrespective of the Chern number of the third moiré band, which sensitively depends on model parameters due to the large energy overlap of this band with remote energy bands. Figure 1 (a) shows the band structure of the Wannier model (dashed cyan lines) compared to the continuum model (black lines) at a displacement field $E_{z}=20\,\mathrm{meV}$. By adding the overlap of each Wannier function with the continuum model’s Bloch states as color, we demonstrate how states from the fourth and fifth bands are essential for a three-orbital description with an overall Chern number sequence $(\nu,\nu,-2\nu)$. The residual spectral weight in remote energy bands is required to render the entire Wannier band manifold topologically trivial. In spite of this, the topmost band of the continuum model is perfectly reproduced in the Wannier model as shown in Fig. 1 (a).
In accordance with previous studies, the topmost band has significant weight in the $T$-orbital Wang et al. (2020); Pan et al. (2020); Zang et al. (2022); Kiese et al. (2022); Klebl et al. (2023); Ryee and Wehling (2023); Zegrodnik and Biborski (2023); Motruk et al. (2023); Wang et al. (2023), but the nontrivial topology invalidates a description as a one-band model. This contrasts MoTe2 bilayers Crépel and Millis (2024b); Qiu et al. (2023) in which more weight lies in the $H_{1,2}$ orbitals, leading to different physical behaviors. In Fig. 1 (b), we display the real-space structure of the three Wannier orbitals. Their exponential localization implies short-ranged hopping parameters as demonstrated in Fig. 1 (a): Disregarding hopping parameters for distances $d\gtrsim 4.5\,|\bm{a}_{M}|$ (with $\bm{a}_{M}$ the moiré lattice vectors) does not change the resulting band structure significantly. A detailed analysis of the energetic and topological properties of the Wannier functions is provided in the SM SM . We carry out the Wannierization for several values of displacement field $E_{z}$ and interpolate the hopping parameters in order to obtain a Hamiltonian for arbitrary $E_{z}$. The resulting density of states is a smooth function of $E_{z}$ [see Fig. 1 (c)], and displays a clear maximum when the Fermi energy meets the VHS of the topmost valence band Schrade and Fu (2021). Along this line, an exceptional point occurs at $E^{c}_{z}\approx 33\,\mathrm{meV}$ (blue circle), where three van-Hove points merge with each other and form a HOVHS Bi and Fu (2021). For $E_{z}<E_{z}^{c}$, the three van-Hove points move towards the $M$-point forming small triangular Fermi pockets around $K^{\nu}$ that are exclusively $T$/$H_{2}$-polarized. In contrast, for $E_{z}>E_{z}^{c}$, the three van-Hove points move towards $\Gamma$, and the Fermi pockets around $K^{\nu}$ disappear. The remaining dispersive part of the Fermi surface is entirely $T$/$H_{2}$-polarized for all values of the displacement field. The multi-orbital Wannier model has the symmetry group $\mathcal{G}_{0}=C_{3z}\otimes U_{\nu}(1)\otimes\mathcal{T}$, where $C_{3z}$ is the point group of the moiré unit cell that contains a three-fold rotation around the $z$-axis, the $U_{\nu}(1)$-symmetry acts in spin-valley space and $\mathcal{T}$ denotes time-reversal symmetry. Due to the absence of cubic terms to lowest order in the continuum description Crépel and Millis (2024b), the system we study theoretically further possesses a mirror-symmetry in each valley $\mathcal{M}_{x}$ that is weakly broken in the physical bilayer but is visible in the fermiology of a single spin-valley sector in Fig. 1 (c).
With the tight-binding Wannier model for arbitrary $E_{z}$ as well as interaction parameters from projection of the dual-gated Coulomb interaction onto the Wannier functions (see SM SM ) at hand, we can formulate the interacting low-energy Hamiltonian as

### v2

@@ Abstract
The recent observation of superconductivity in the vicinity of Fermi surface reconstructed insulating or metallic states has established twisted bilayers of WSe2 as an exciting platform to study the interplay of strong electron-electron interactions, broken symmetries and topology. In this work, we use a first-principles, material-specific theoretical treatment that is unbiased with respect to electronic instabilities to study the emergence of electronic ordering in twisted WSe2 driven by gate-screened Coulomb interactions. We construct exponentially localized moiré Wannier orbitals that faithfully capture the bandstructure and topology of the system, project the gate-screened Coulomb interaction onto them and use unbiased functional renormalization group techniques to resolve the momentum and orbital structure of the leading instabilities and the relevant energy scales. We find an interplay between intervalley-coherent antiferromagnetic (IVC-AFM) order and chiral, mixed-parity $d/p$-wave superconductivity for carrier concentrations near a displacement field and twist-angle-tunable van-Hove singularity. Our microscopic approach establishes incommensurate IVC-AFM spin fluctuations as the dominant electronic mechanism driving the formation of superconductivity in $\theta=5.08^{\circ}$ twisted WSe2 and explains key aspects of recent experiments including the asymmetric density dependence of the spin ordering with respect to the van-Hove line, the single and double-peak structure of the DOS in the ordered (hole-doped) IVC-AFM phase, the emergence of superconductivity as the density is varied across the van-Hove line and the evolution of the displacement field-density phase diagram with twist angles between $3.7^{\circ}\dots 5^{\circ}$. We show that the interplay of electronic correlations and non-trivial quantum geometry in tWSe2 manifests in orbital-selective order parameters associated to the IVC-AFM and SC states which are detectable by local spectroscopy measurements.
@@ I Introduction
Transition metal dichalcogenide (TMD) heterostructures have arisen as fertile and highly versatile platforms for exploring the interplay between topology and strong electronic correlations [1], hosting correlated insulators [2, 3, 4], spin-ordered metals [5, 6, 7], rich multi-orbital physics [8, 9], Feschbach resonances [10, 11, 12, 13, 14, 15, 16], quantum critical behavior [17, 18], excitonic insulators [19, 20, 21, 22], topological Kondo physics [23, 24, 25, 26, 27, 28], quantum anomalous and quantum spin-Hall insulators [29, 30], ferromagnetic fractional quantum anomalous phases [31, 32, 33, 34], or gapped spin liquids [35, 36]. Additionally, early experiments on $5.0^{\circ}$ twisted WSe2 hinted at the existence of superconducting phases in moiré TMDs [2]; a finding recently confirmed by sub-Kelvin transport measurements on $3.65^{\circ}$ [37], $4.6^{\circ}$ [38] and $5.1^{\circ}$ [39] twisted samples. The discovery of superconductivity in these highly tunable moiré TMDs has sparked many theoretical studies [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 48] aimed at understanding the origin and properties of strongly correlated topological superconductivity. Exciting proposals for trapping non-abelian anyons at clean, gate-defined interfaces have also appeared [51, 52, 53, 54].
@@
The association of superconductivity with proximity to the non-integer carrier concentrations at which the Fermi surface contains a van-Hove singularity (VHS), and the proximity to an apparent metallic antiferromagnetic phase suggests a weak-to-moderate coupling scenario in the $5.1^{\circ}$ device, where the effects of electron-electron interactions are amplified by the density of states divergence associated with the van-Hove point. The association of superconductivity at particular values of the displacement field may then be related to proximity to a higher order van-Hove singularity (HOVHS) with a stronger density of states divergence appearing as the band theory is tuned [55]. Recent literature has presented various versions of this scenario. Interactions considered include spin-fluctuations [42, 40, 48], over-screened Coulomb interactions [46] and phenomenological attraction [44, 47]. While these treatments provide important insights, they are based on particular ansätze and come to different conclusions regarding the symmetry of the superconducting state.
@@
Moreover, theory should take into account the non-trivial topological properties [56, 57, 58] of the low-lying hole bands. In addition to potentially endowing the superconducting state with interesting properties [46], they obstruct an effective single-orbital description of the topmost moiré valence band and imply that interactions projected onto the top-most band generically have a long-ranged structure. A minimal (topologically unobstructed) faithful low-energy model of tWSe2 comprises three orbitals per valley that are centered on the triangular and honeycomb sites of the moiré superlattice [57, 59]. Such a model accurately accounts for the topological entanglement of the topmost valence bands with Bloch states far below the Fermi energy and for the subtle aspects of the fermiology and its dependence on the displacement field. In this multi-orbital Wannier basis, the gate-screened Coulomb interactions relevant for the experimental setup can be cast into fairly local interactions, which ensures the applicability of established many-body techniques beyond the mean-field level [47, 46], including the functional renormalization group [60, 61, 62, 63, 64] (FRG), dynamical mean-field theory [65], or tensor network based approaches [66, 67]. A similar situation arises in twisted bilayers of graphene, where a multi-band “heavy fermion” description is argued to best represent the physics [68, 69].
@@
In this work, we go beyond previous treatments of correlated phases in tWSe2 and provide a first-principles, theoretically unbiased analysis of correlated states in the weak-to-moderate interaction regime. Our theory is based on physically realistic repulsive gate-screened Coulomb interactions projected onto a multi-band model based on exponentially localized Wannier functions that capture the quantum geometry and the intricacies of the band structure. Using functional renormalization group techniques that are unbiased in the sense of not a priori favoring any particular electronic instability, we map out the interacting phase diagram of tWSe2 in the twist angle range $\theta=3.7^{\circ}\dots 5^{\circ}$ from first principles. We demonstrate that intervalley-coherent anti-ferromagnetic (IVC-AFM) order and extended regions of chiral $d/p$-wave superconductivity coexist along the gate-tunable van-Hove line. The unbiased characterization of particle-hole and particle-particle instabilities as well as their momentum-agnostic properties in conjunction with the subtle fermiology near the van-Hove line whose analysis is enabled by our choice of methods permits us to show that the electronic driving mechanism for superconductivity in tWSe2 is primarily facilitated by spin fluctuations around regions of incommensurate IVC-AFM order below the higher order van-Hove singularity, where the absence of perfect nesting strengthens particle-particle fluctuations. The topology of the system manifests in orbital-selective order parameters that have weight only in certain domains of the tWSe2 superlattice. The interplay of electronic correlations and the non-trivial quantum geometry of the moiré flat bands manifests in orbital-selective order parameters that have weight only in certain domains of the tWSe2 superlattice. Our microscopic approach can explain the most pivotal features observed in recent experimental measurements. For the the twist angle of $\theta=5.08^{\circ}$ this includes the unusual density dependence of the particle-hole instabilities with respect to the van-Hove line, the transition from a single-peak to a double-peak structure in the density of states (DOS) as well as the subtle interplay of IVC-AFM order and superconductivity along the tunable van-Hove line. As the twist angle is decreased below $\theta=5.08^{\circ}$, our theory predicts a successive shift of IVC-AFM domains towards the commensurate filling $n/n_{0}\sim-1$ and smaller values of the displacement field accompanied by an increasing asymmetry of particle-hole instabilities with respect to the van-Hove line. The superconducting domains in the phase diagram are eventually displaced from the van-Hove line. The accompanying decrease in the density of states is found to be more important than the increase in interaction, leading to a reduction in critical temperature. Our theoretical analysis indicates that the leading SC order parameter is of mixed-parity, chiral $d/p$-wave character, for all twist angles studied, however, its orbital decomposition changes such that the system transitions from a multi-orbital to a single-orbital superconductor.
This paper is structured as follows: In Section II we present the basic physics of the system. In Section III.3, we describe how to construct a faithful, multi-orbital Wannier model that captures the intricate fermiology and quantum geometry of tWSe2 and accounts for the dual-gated Coulomb interaction between the electrons and discuss our treatment of electronic correlations within the framework of the functional renormalization group (FRG) that allows an unbiased, beyond mean-field characterization of emergent phases in tWSe2. In Section III.3, we map out the FRG phase diagram of $\theta=5.08^{\circ}$ and unravel a delicate interplay between IVC-AFM order and chiral $d/p$-wave superconductivity along the displacement-field tunable van-Hove line. In Section V, we first provide a detailed characterization of the IVC-AFM state and access quasiparticle properties by performing a FRG+MF analysis in the ordered phase. In particular, we demonstrate how significant electronic screening contributions accumulated by particle-particle scattering within FRG leads to a suppression of effective Hubbard-like interactions, which bridges to the observed over-estimation of electronic order observed in mean-field (Hartree-Fock) approaches. In Section VI, we discuss the electronic mechanism leading to the formation of chiral $d/p$-wave superconductivity in tWSe2 and its relation to the nearby IVC-AFM domains in the phase diagram. We provide an in-depth analysis of spectral properties in the SC phase and unravel a strong orbital-selectivity that manifests in nodal signatures in the local density of states (LDOS) and is related to the topological properties of the Wannier model. In Section VII, we provide a detailed analysis on the role of long-ranged density-density interactions in the formation of IVC-AFM and SC order. Having established our methodology for $\theta=5.08^{\circ}$ tWSe2, we finally map out the twist angle evolution of IVC-AFM and SC order in the range of $\theta=3.7\ldots 5^{\circ}$ in Section VIII.
@@
Figure 1: Fermiology and Wannierization of $\theta=5.08^{\circ}$ tWSe2 for an external displacement field $E_{z}=20\,\mathrm{meV}$. (a) Continuum bandstructure in valley $\nu=+1$ (black) and spectral overlap of its Bloch states with a three-orbital model comprising Wannier orbitals centered at the MM stacking regions ($T$, $s$-orbital) and XM/MX stacking regions ($H_{1}$, $p_{+}$-orbital)/($H_{2}$, $p_{+}$-orbital) of the moiré superlattice (tri-colored map and inset). The bandstructure of the effective Wannier model truncated after 4.5 moiré lattice vectors is indicated by the green dashed line. (b) Amplitude $|\psi_{l}\bm{(}r)|^{2}$ of the Wannier functions ($T,H_{1},H_{2}$) on the two layers ($l=0,1$) that capture the spectral weight of the continuum model. (c) Density of states as function of holes per moiré unit cell $n/n_{0}$ and displacement field $E_{z}$. The upper panel shows snapshots of the Fermi surface along the van-Hove line and its orbital polarization using the same colormap as in (a) for valley $\nu=+1$. The Fermi surface of the opposite valley $\nu=-1$ is shown in gray. The green dashed line indicates the mini-Brillouin zone of tWSe2. The blue circle indicates the position of the higher-order VHS, where three van-Hove points merge at $K^{\nu}$ as demonstrated in the insets. The three van-Hove points are moved towards $\Gamma$ ($M$) for increasing (decreasing) value of the external displacement field along the van-Hove line as indicated by the blue triangle (square).
II WSe2 overview
Twisted WSe2 (tWSe2) consists of two monolayers of WSe2 stacked at a relative twist angle $\theta$ and encapsulated in a structure including top and bottom gates. The twist, in combination with the electrostatic potential from the local arrangement of WSe2 bilayers and the interlayer electronic hybridization, produces a large unit cell with two dimensional “moiré” bands [70, 56]. The average of the potentials on the top and bottom gates tunes the electronic density over wide ranges. The difference in voltage between the top and bottom gates introduces an interlayer potential difference, or “displacement field”, which shifts electronic wave functions and charge density between the layers, modifying the electronic band structure. At zero and small displacement fields and hole concentrations not too different from one per unit cell, the bands near the Fermi surface are derived from a complicated combination of states distributed over both layers and originating from various regions of the moiré unit cell. This complicated combination of states leads to a non-trivial Fermi surface shown in Fig. 1(b) and non-trivial quantum geometry that (as will be seen) leads to a nontrivial Chern number of the superconducting state. Increasing the displacement field at fixed carrier density drives the system into a "layer-polarized" regime in which the near Fermi surface states are drawn primarily from one layer and one region of the moiré unit cell. In this regime, superconductivity is absent and the quantum geometry of the bands is simple. At a fixed twist angle, the system’s emergent behavior can be explored in a phase diagram defined by electronic density and displacement field. As the density varies at a fixed displacement field, the Fermi surface can pass through a van-Hove singularity (VHS), where the density of states (DOS) diverges. Of particular relevance for this study is that both the density at which a van-Hove singularity occurs and the strength of the resulting divergence in the DOS depend on twist angle and are tunable over wide ranges by varying the displacement field.
In the $5.1^{\circ}$ tWSe2 device [39], the superconducting state is observed near hole densities $n/n_{0}\sim n_{\text{VHS}}\neq-1$ at which the Fermi surface passes through the van-Hove singularity, and only for a small range of displacement fields. The superconducting phase is adjacent to a metallic region characterized by an increase in longitudinal resistance and a very rapid non-monotonic doping dependence of the Hall resistivity. This unusual metallic region has been interpreted [39] as arising from the emergence of antiferromagnetic (AFM) spin ordering that partially gaps out the Fermi surface. The densities at which superconductivity and antiferromagnetism occur are not close to one hole per moiré unit cell and the experimentally important displacement fields are significant, on the order of half of the field required to enter the layer-polarized regime. As the twist angle is decreased, the density range where superconductivity and antiferromagnetism are observed moves closer to one hole per unit cell and the relevant displacement field decreases. In the $3.65^{\circ}$ device [37], the antiferromagnetic phase is centered at hole densities of $n/n_{0}=-1$ per moiré unit cell and is insulating, presumably because at this commensurate carrier concentration the commensurate antiferromagnetic order can fully gap the Fermi surface. The antiferromagnetic insulating phase exists for displacement fields between a small but non-zero value, and the “layer-polarization" displacement field $E_{Z}^{\text{lp}}$ beyond which the system enters the layer polarization regime. In this device superconductivity is also found only for carrier concentrations near $n/n_{0}=-1$ per moiré unit cell and small displacement fields adjacent to the insulating state.
In the next Section we present our theoretical model, which encodes the twist angle-dependent fermiology and the relevant interacting physics.
III Model and method
III.1 Multi-orbital Wannier model
@@
The non-interacting electronic structure of hole-doped tWSe2 is accurately described by a continuum model in which the spin-valley locked holes of both monolayers are treated within the effective mass approximation, and are coupled by moiré potentials that slowly vary on the atomic scale [70]. Due to the emergent spin-valley locking in monolayer WSe2, the low-energy states can be assigned a well-defined valley index $\nu=\pm 1$ and states from opposite spin-valley sectors are related by virtue of time-reversal symmetry $\mathcal{T}$. The effects of out-of-plane electric fields are captured by a layer-dependent potential energy $\pm E_{z}$. For twist angles $\theta\gtrsim 3^{\circ}$, in-plane lattice relaxation is weak, and the associated moiré fields are characterized by only a few harmonics [71, 72, 73, 74, 75]. Here, we use the continuum model from Ref. [56], whose coefficients have been fitted to large-scale ab-initio calculations, see Supplementary Materials (SM) [76] for details.
@@
As most many-body methods require a symmetry-preserving lattice description of the system under scrutiny, we seek to construct a Wannier representation of the topmost valence band states in each valley based on the three-orbital model put forward in Refs. [57, 59]. This model is based on generic symmetry arguments and comprises one triangular lattice site $T$ in the MM stacking regions of the triangular superlattice (Wyckoff position 1a), which has weight on both layers, as well as two honeycomb lattice sites $H_{1,2}$ in the XM/MX stacking regions (Wyckoff positions 2a,b) with dominant weight on either one of the layers. At a twist angle of $\theta=5.08^{\circ}$, however, the Chern number sequence of the first three valence bands of tWSe2 in valley $\nu=\pm 1$ is $(\nu,\nu,\nu)$ for the parameters of Ref. [56], prohibiting an exponential localization of Wannier functions centered at aforementioned Wyckoff positions constructed from the three topmost bands alone. We resolve this issue by employing a single-shot Wannierization scheme [77, 78, 79] in which Bloch states from the entire valence band manifold are considered for the projection and a proper subspace selection is instead achieved by weighting Bloch states according to their energetic hierarchy and their weight around the desired orbital center. Compared to the layer-polarization scheme presented in Ref. [57], the energetic subspace selection ensures to construct exponentially localized Wannier functions irrespective of the Chern number of the third moiré band, which sensitively depends on model parameters due to the large energy overlap of this band with remote energy bands. Figure˜1 (a) shows the band structure of the Wannier model (dashed cyan lines) compared to the continuum model (black lines) at a displacement field $E_{z}=20\,\mathrm{meV}$. By adding the overlap of each Wannier function with the continuum model’s Bloch states as color, we demonstrate how states from the fourth and fifth bands are essential for a three-orbital description with an overall Chern number sequence $(\nu,\nu,-2\nu)$. The residual spectral weight in remote energy bands is required to render the entire Wannier band manifold topologically trivial. In spite of this, the topmost band of the continuum model is perfectly reproduced in the Wannier model as shown in Fig.˜1 (a).